# captionUnknown document class (or package)

### v1

@@ Abstract
Inter-agent relative localization is critical for many multi-robot systems operating in the absence of external positioning infrastructure or prior environmental knowledge. We propose a novel inter-agent relative 3D pose estimation system where each participating agent is equipped with several ultra-wideband (UWB) ranging tags. Prior work typically supplements noisy UWB range measurements with additional continuously transmitted data, such as odometry, leading to potential scaling issues with increased team size and/or decreased communication network capability. By equipping each agent with multiple UWB antennas, our approach addresses these concerns by using only locally collected UWB range measurements, a priori state constraints, and detections of when said constraints are violated. Leveraging our learned mean ranging bias correction, we gain a 19% positional error improvement giving us experimental mean absolute position and heading errors of 0.24m and $\bm{{}^{\circ}}$ respectively. When compared to other state-of-the-art approaches, our work demonstrates improved performance over similar systems, while remaining competitive with methods that have significantly higher communication costs. Additionally, we make our datasets available.
@@ I Introduction
Multi-robot systems can be used to improve the efficiency and robustness of large-scale tasks such as search & rescue [1], warehouse automation [2], and planetary exploration [3]. To operate and parallelize effectively, these systems typically need to know where the agent (and its peers) are located in a common reference frame. In practice this is often achieved by localizing within an a priori map or using an external measurement system like GPS or motion capture (mocap) [4, 5]. If these technologies are unavailable or infeasible, common approaches utilize both relative localization [6] and multi-agent SLAM [7] techniques.
@@
Figure 1: Diagram of proposed system. Here three agents fly at different altitudes while performing real-time 3D relative pose estimation. Each agent is equipped with six ultra-wideband (UWB) antennas, each capable of performing pairwise relative ranging between all other agents’ individual antennas. By using trilateration, an improved sensor model, and a priori state constraints about altitude/roll/pitch, agents can perform instantaneous estimation entirely with locally collected UWB measurements (i.e., without the need to continuously transmit other measurements, such as odometry). Additionally, each agent locally monitors its a priori constraints via downward facing LiDAR and IMU, and thus only needs to transmit one-off messages with the swarm if these assumptions change or are violated.
@@
Within the last decade, ultra-wideband (UWB) has matured into a reliable, inexpensive, and commercially available RF solution for data transmission, relative ranging, and localization – UWB now comes as standard issue in many popular smartphone devices [8]. For robotics, UWB has several properties of note: precision of approximately 10cm, ranges up to 100m, resilience to multipath, operates in non-line of sight (NLOS) conditions, low power consumption, and 100Mbit/s communication speeds [9]. Recent devices even extend the recommended and operational ranges to 300m and 500m respectively [10]. Nevertheless, UWB measurements are not immune from ranging errors or noise (see Section III-A), the modeling and correction of which is an active area of research [11, 12, 13, 14].
@@
A common approach in UWB relative localization work fuses noisy UWB ranging measurements with additional continuously transmitted data, such as odometry [15] and visual inter-agent tracks [16, 17]. While these approaches achieve low absolute position error (APE) and absolute heading error (AHE), there are two prevalent shortcomings: (1) They often use a simplistic UWB measurement noise model (i.e., zero mean Gaussian), which then requires the use of supplementary measurements to compensate. (2) Reliance on these supplementary measurements (often not locally111locally” refers to a measurement collected onboard, and thus can be used to estimate other agents’ pose without needing to be transmitted. collected, e.g., odometry), mandates their continuous transmission between agents, potentially impacting scalability to increased swarm size or decreased communication throughput.
@@
Our previous work [18] used UWB to demonstrate an instantaneous222instantaneously” implies the target’s position is fully determined (i.e., observable) using only current measurements. As such, this approach does not require a dynamics model or measurements over a period of time. multi-tag approach to relative 2D pose estimation that achieved superior mean position accuracy and competitive performance on other metrics to Cao et al. [15] (the most comparable state-of-the-art work) using only local UWB measurements. These results were achieved by a trilateration nonlinear least squares (NLLS) optimization problem that, leveraging an improved UWB sensor model, accounted for various sources of measurement error.
@@
This letter extends our prior 2D work [18] to an instantaneous multi-tag relative pose estimate in 3D environments (Figure 1). Similar to our previous work, a key objective is to develop an approach that minimizes the communication load (i.e., we are willing to trade off absolute accuracy for reduced communication overhead). In 3D environments, each agent’s altitude/roll/pitch are locally measurable with respect to the world frame (e.g., via altimeter or downward facing LiDAR and IMU). By constraining each agent to an altitude/roll/pitch envelope, this information only needs to be transmitted once a priori. These assumptions can then be locally monitored by each agent, with one-off333one-off” denotes communication is intended to be infrequent or non-existent (i.e., not routine or continuous). Specifically, messages exist only to notify an a priori constraint has been violated or changed. communication occurring only to alert the swarm a constraint has been violated (i.e., one-off messages only communicate a new constraint envelope if it has changed, or let the swarm know to ignore the agent until further notice). By leveraging and monitoring these state constraints, we simplify the 6-DoF optimization (i.e., 3D position and orientation) to a 3-DoF optimization (i.e., $x$, $y$, yaw with known $z$, roll, pitch) that, with fewer free variables, more reliably produces 3D localization solutions. Thus our approach provides a 3D relative pose using only locally collected UWB range measurements, a priori state constraints, and detections of constraint violations.
@@
This paper’s contributions are: (1) An in-depth analysis and modeling of the observed noise characteristics of 3D UWB ranging measurements (Section III). (2) A 3D solution for UWB-based relative localization, formulated as a robust nonlinear least squares optimization with minimal communication footprint and locally monitored assumptions. We utilize a learned measurement bias correction as a function of relative elevation that accounts for observed noise properties, improving our mean absolute position error by 19% (Section V). (3) Experimental results that demonstrate the merits of our proposed solution, a mean APE and AHE of 0.24m and ${}^{\circ}$ respectively without continuously transmitting measurements (Section VI). (4) Our datasets, over 200 hours of pairwise UWB range measurements with ground truth, for use by the larger robotics community.444 https://github.com/mit-acl/murp-datasets
@@
Figure 2: Plots demonstrating the UWB noise and geometry characteristics outlined in Section III. (a) A histogram of our entire dataset’s UWB measurements binned by range error. Demonstrates a non-zero mean and a long tail (i.e., violates the zero mean Gaussian assumption used by many works). (b) The same data plotted as error with respect to relative elevation. Here, we see the measurement error’s mean and variance significantly change with relative elevation. The dotted blue line represents a learned 6-degree polynomial fit of measurement bias. (c) Simple example demonstrating dilution of precision (DOP) in a 2D ranging scenario. Here, the red and green points have known baselines of 3m and 6m respectively. Each red/green point independently measures the distance to the blue point with a consistent uncertainty of ±$\pm 1$m. To estimate the position of the blue point, the red and green points each intersect their uncertain measurements, respectively creating red and green feasibility regions. Despite having the same uncertainty on all range measurements, the green points (i.e., the larger baseline) produce a smaller uncertainty region than the red points. (d) This diagram demonstrates how three (or more) ranging antennas within a single plane produce a pair of ambiguous solutions (i.e., if all antennas are in the base’s plane $z=0$, while the target’s $x$ and $y$ coordinates are fully observable, the target’s altitude has an ambiguity between ±$\pm z$).
@@ II Related Works
2D
Guo et al.     (2020) [19]
Single Bounded Error Assumption UWB + Velocity
Cao et al.     (2021) [15]
Many (4) NLLS UWB + Odometry
Zheng et al.     (2022) [20]
Many (2 or 4) ZM Gaussian + NLOS Rejection UWB + VIO
Zhang et al.     (2023) [21]
Many (2 or 4) ZM Gaussian + Outlier Rejection UWB + Point Clouds + Odom + Keyframes
[Our Prev. Work]     (2022) [18]
Many (4) NLLS + Data Informed Sensor Model UWB
3D
Xu et al.     (2020) [16]
Single ZM Gaussian UWB + VIO + Visual Tracks
Qi et al.       (2020) [22]
Single ZM Gaussian UWB + IMU + GPS Heading
Xianjia et al.     (2021) [23]
Many (2 or 4) NLLS UWB
Xu et al.     (2022) [17]
Single ZM Gaussian + Occlusion Rejection UWB + VIO + Visual Tracks
Xun et al.     (2023) [24]
Single ZM Gaussian UWB + IMU + IR Visual Tracks
[This Work]     (2023)
Many (4 or 6) NLLS + Data Informed Sensor Model UWB + One-Off Assumption Violations
@@
Zhang et al. [21] presents a centralized 2D range-aided cooperative localization and consistent reconstruction system, merging a tightly coupled visual odometry and multi-tag ranging front-end with a pose graph optimization (PGO) back-end. This work differs from ours by being 2D and relying on continuous transmission of point cloud, odometry, and keyframe information. Omni-swarm [17] is a decentralized 3D swarm estimation system. This work differs from ours by relying primarily on visual tracks of neighboring agents from each agent’s omnidirectional camera – omitting UWB measurements only degrades estimation by approximately 0.01m. Additionally, each agent has only a single UWB tag and requires continuously sharing measurements between all pairs of agents, scaling poorly with larger swarms. CREPES [24] presents a custom compact hardware module, which tightly couples an IMU, UWB, IR LEDs, and IR fish-eye camera. The measurements are fused into 6-DoF relative pose via a centralized error-state Kalman filter (ESKF) and PGO, while also having the ability to provide an instantaneous estimate from a single frame of sensor measurements. This work differs from ours by relying heavily on visual tracks over UWB (similar to [17]). Additionally, the system relies on continuously transmitted measurements. Thus the state of the art approaches in [21, 17, 24] achieve APE and AHE on the order of 0.10 m and $1^{\circ}$, respectively, but they require additional sensors (e.g., LiDAR, fish-eye cameras, etc.) that must continuously transmit measurements, making them scale poorly to scenarios with larger swarm sizes or reduced communication throughput.
@@ III Extensions to 3D Environments
The following subsections motivate our hardware and algorithmic choices for operating in 3D (Section III-C). We first characterize the UWB ranging error of the Nooploop LinkTrack P-B [10] from a set of initial experiments [^5] While these effects should generalize to other UWB ranging devices, given the “black box” nature of commercial sensors, exact manifestations may differ between devices. However, most works in Table I use a version of the Nooploop LinkTrack P [10] for UWB ranging, except for [19, 23], which use the PulsON 440 and an unspecified device respectively. Thus, we expect similar noise characteristics among most related works. (Section III-A), and then provide the observability and error properties of trilateration-based localization (Section III-B).
@@ III-A Characterizing UWB Noise
Long Tail Distributions: Contrary to common noise assumptions in Table I, individual UWB ranging errors appear neither zero mean nor Gaussian. Instead, error distributions have long positive tails (see Figure 1(a)). A trend towards positive bias can be attributed to the many ways positive ranging error can be introduced to a UWB measurement (e.g., change of propagation medium, multipath). Furthermore, installing SMA cables between the RF device and antenna adds a consistent positive bias to all measurements.
@@ III-B Geometry of Ranging-based Pose Estimation
Dilution of Precision: Most commonly associated with GPS, dilution of precision (DOP) is a trilateration sensitivity analysis for quantifying how 1D ranging errors propagate to 2D/3D point estimation error [31]. Specifically, given trilateration’s nonlinear nature, even identically noisy 1D range measurements can produce wildly different 2D/3D positional uncertainties – this is entirely determined by the given base station geometry (see Figure 1(c)). In general we note that increasing the distance between base stations (i.e., the “baseline”), or being closer to the base stations, reduces trilateration’s sensitivity. See [20, 21] for a detailed discussion of DOP as it pertains to 2D UWB localization. This also accounts for the improved accuracy seen by [23] for configurations with larger baselines. Thus, the accuracy of trilateration systems with vastly different baselines or operating ranges can be misleading.
@@ III-C System Design
Based on the noise characteristics (Section III-A) and underlying geometry (Section III-B), we propose the system design shown in Figure 1 for our multirotor drones. Specifically, each agent is equipped with 6 ranging antennas each attached at the end of their propeller arms (i.e., 0.31m from the center) to maximize baseline distances and create better DOP geometry. Since there are 6 coplanar antennas, this is an over-constrained problem that still has a ±$\pm z$ ambiguity across the $z=0$ plane. Although an additional nonplanar antenna can be to each agent, it is challenging to get a significant $z$-baseline without creating poor flight conditions (i.e., prohibitively heavy or awkward center of mass). Instead, we note that altitude/roll/pitch can be measured within the flat-ground world frame locally via altimeter or downward facing LiDAR and onboard IMU. In many drone applications, it is common for multirotors to maintain roll/pitch near $0^{\circ}$ and a constant altitude. Thus we can specify a minimalist communication protocol that shares these intended constraints a priori and then monitors them locally to ensure they are maintained. One-off communication occurs only when a constraint is locally detected to have been violated/changed (see Figure 2(b)). Additionally, having a constant $z$ constraint helps mitigate the increased variance associated with relative elevation (see Figure 1(b)).
@@
Figure 3: (a) Annotated diagram of three UGV agents used in Section VI-B. UGV agents are designed with the same baseline as the UAV in Figure 1, making them comparable surrogates to the UAV and its experiments in Section VI-A. (b) Diagram demonstrating agent $B$ locally detecting violations in its a priori constraints, triggering a one-off communication.
IV Preliminaries
IV-A Pose Parameterization
Given a set of 3D reference frames $\mathcal{F}$, consider any pair of frames $A\in\mathcal{F}$ and $B\in\mathcal{F}$. Let $\mathbf{T}^{A}_{B}\in SE(3)$ be the relative pose between frames $A$ and $B$ (i.e., transforms a point from $B$’s frame into $A$’s frame). We note $\mathbf{T}^{A}_{B}$ is a 6-DoF value that can be equivalently parameterized as the 2-tuple $\langle\mathbf{R}^{A}_{B},\mathbf{t}^{A}_{B}\rangle$ or 6-tuple $\langle x^{A}_{B},y^{A}_{B},z^{A}_{B},\alpha^{A}_{B},\beta^{A}_{B},\gamma^{A}_{B}\rangle$, where:
  • $\mathbf{R}^{A}_{B}\in SO(3)$ is the relative rotation between $A$ and $B$ parameterized by the 3-tuple of Euler angles $\langle\alpha,\beta,\gamma\rangle$, where $\alpha^{A}_{B}\in[-180^{\circ},180^{\circ}]$ is relative roll, $\beta^{A}_{B}\in[-90^{\circ},90^{\circ}]$ is relative pitch, and $\gamma^{A}_{B}\in[-180^{\circ},180^{\circ}]$ is relative yaw. The use of Euler angles over quaternions is motivated by the direct roll/pitch constraints to be introduced Equation 7. We define $\mathbf{R}(\alpha,\beta,\gamma)$ as the $\mathbb{R}^{3\times 3}$ rotation matrix with respect to $\mathbf{R}_{x}(\alpha)$, $\mathbf{R}_{y}(\beta)$, $\mathbf{R}_{z}(\gamma)$, the $\mathbb{R}^{3\times 3}$ rotation matrices about the $x$, $y$, $z$ axes respectively. Thus $\mathbf{R}^{A}_{B}$ can be equivalently represented as a $\mathbb{R}^{3\times 3}$ rotation matrix in the following ways:
    $\mathbf{R}^{A}_{B}\triangleq\mathbf{R}(\alpha^{A}_{B},\beta^{A}_{B},\gamma^{A}_{B})\triangleq\mathbf{R}_{z}(\gamma^{A}_{B})\mathbf{R}_{y}(\beta^{A}_{B})\mathbf{R}_{x}(\alpha^{A}_{B})$
  • $\mathbf{t}^{A}_{B}\in\mathbb{R}^{3}$ is a relative translation vector between $A$ and $B$ parameterized by the 3-tuple $\langle x^{A}_{B},y^{A}_{B},z^{A}_{B}\rangle$. Thus, $\mathbf{t}^{A}_{B}$ can be equivalently represent in the following ways:
    $\mathbf{t}^{A}_{B}\triangleq\mathbf{t}(x^{A}_{B},y^{A}_{B},z^{A}_{B})\triangleq\begin{bmatrix}x^{A}_{B}&y^{A}_{B}&z^{A}_{B}\end{bmatrix}^{\top}$
IV-B Rigid Body Transformations
For any frame $F\in\mathcal{F}$, let $\mathbf{p}^{F}\in\mathbb{R}^{3}$ be a point in frame $F$. Given $\mathbf{T}^{A}_{B}$ (i.e., the relative pose between frames $A$ and $B$), points $\mathbf{p}^{A}$ and $\mathbf{p}^{B}$ (i.e., the same point but in $A$’s or $B$’s frame respectively) are related by:
$\mathbf{p}^{A}=\mathbf{R}^{A}_{B}\mathbf{p}^{B}+\mathbf{t}^{A}_{B}$
which can equivalently and succinctly be written in homogeneous coordinate notation [32] as:
$\breve{\mathbf{p}}^{A}=\mathbf{T}^{A}_{B}\breve{\mathbf{p}}^{B}$
where $\mathbf{T}^{A}_{B}\in\mathbb{R}^{4\times 4}$ and $\breve{\mathbf{p}}^{F}\in\mathbb{R}^{4}$ are defined as:
$\mathbf{T}^{A}_{B}\triangleq\begin{bmatrix}\mathbf{R}^{A}_{B}&\mathbf{t}^{A}_{B}\\ \mathbf{0}^{\top}&1\end{bmatrix}\qquad\qquad\breve{\mathbf{p}}^{F}\triangleq\begin{bmatrix}\mathbf{p}^{F}\\ 1\end{bmatrix}$
Thus $\mathbf{T}^{A}_{B}$ is equivalently represented in the following ways:
$\mathbf{T}^{A}_{B}$ $\triangleq\mathbf{T}(x^{A}_{B},y^{A}_{B},z^{A}_{B},\alpha^{A}_{B},\beta^{A}_{B},\gamma^{A}_{B})$
$\triangleq\begin{bmatrix}\mathbf{R}^{A}_{B}&\mathbf{t}^{A}_{B}\\ \mathbf{0}^{\top}&1\end{bmatrix}\triangleq\begin{bmatrix}\mathbf{R}(\alpha^{A}_{B},\beta^{A}_{B},\gamma^{A}_{B})&\mathbf{t}(x^{A}_{B},y^{A}_{B},z^{A}_{B})\\ \mathbf{0}^{\top}&1\end{bmatrix}$
V Technical Approach

### v3

@@ Abstract
Inter-agent relative localization is critical for many multi-robot systems operating in the absence of external positioning infrastructure or prior environmental knowledge. We propose a novel inter-agent relative 3D pose estimation system where each participating agent is equipped with several ultra-wideband (UWB) ranging tags. Prior work typically supplements noisy UWB range measurements with additional continuously transmitted data (e.g., odometry) leading to potential scaling issues with increased team size and/or decreased communication network capability. By equipping each agent with multiple UWB antennas, our approach addresses these concerns by using only locally collected UWB range measurements, a priori state constraints, and event-based detections of when said constraints are violated. The addition of our learned mean ranging bias correction improves our approach by an additional 19% positional error, and gives us an overall experimental mean absolute position and heading errors of 0.24m and 9.5 respectively. When compared to other state-of-the-art approaches, our work demonstrates improved performance over similar systems, while remaining competitive with methods that have significantly higher communication costs. Additionally, we make our datasets available.
$\footnotesize Data/Video:\url{https://github.com/mit-acl/murp-datasets}$
@@ I Introduction
Figure 1: Diagram of proposed system. Here three agents fly at different altitudes while performing real-time 3D relative pose estimation. Each agent is equipped with six ultra-wideband (UWB) antennas, each capable of performing pairwise relative ranging between all other agents’ individual antennas. By using trilateration, an improved sensor model, and a priori state constraints about altitude/roll/pitch, agents can perform instantaneous estimation entirely with locally collected UWB measurements (i.e., without the need to continuously transmit other measurements, such as odometry). Additionally, each agent locally monitors its a priori constraints via downward facing LiDAR and IMU, enabling an event-based communication model that only transmits if said assumptions change or are violated. The pictured drone is used in the experiments outlined in Section V-A.
@@
Within the last decade, ultra-wideband (UWB) has matured into a reliable, inexpensive, [^1] At time of writing, UWB sensors sell for ~\$10 as an integrated circuit and between \$20-\$100 as a development board or plug-and-play sensor.and commercially available RF solution for data transmission, relative ranging, and localization. For robotics, UWB has several properties of note: precision of 10cm, ranges up to 100m, resilience to multipath, operates in non-line of sight (NLOS) conditions, low power consumption, and 100Mbit/s communication speeds [1]. Recent devices even extend the recommended and operational ranges to 300m and 500m respectively [2]. Nevertheless, UWB measurements are not immune from ranging errors or noise (see Section III-A), the modeling and correction of which is an active area of research [3, 4, 5, 6].
@@
A common approach in UWB relative localization work fuses noisy UWB ranging measurements with additional continuously transmitted data, such as odometry [7] and visual inter-agent tracks [8, 9]. While these approaches achieve low absolute position error (APE) and absolute heading error (AHE), there are two prevalent shortcomings: (1) They often use a simplistic UWB measurement noise model (i.e., zero mean Gaussian), which then requires the use of supplementary measurements to compensate. (2) They rely on these supplementary measurements (often not locally222locally” refers to a measurement collected onboard, and thus can be used to estimate other agents’ pose without needing to be transmitted. collected, e.g., odometry), which mandates their continuous transmission between agents and impacts system scalability with increased swarm size or decreased communication throughput.
@@
Our previous work [10] used UWB to demonstrate an instantaneous333instantaneous” implies the target’s position is fully determined (i.e., observable) using only current measurements. As such, this approach does not require a dynamics model or measurements over a period of time. Furthermore, since estimation does not rely on integrating velocity or acceleration, the estimation should not drift/diverge over time. multi-tag approach to relative 2D pose estimation that achieved superior mean position accuracy and competitive performance on other metrics to Cao et al. [7] (the most comparable state-of-the-art work) using only local UWB measurements. These results were achieved by a trilateration nonlinear least squares (NLLS) optimization problem that, leveraging an improved UWB sensor model, accounted for various sources of measurement error.
@@
This letter extends our prior 2D work [10] to an instantaneous multi-tag relative pose estimate in 3D environments (Figure 1). Similar to our previous work, a key objective is to develop an approach that minimizes the communication load (i.e., we are willing to trade absolute accuracy for reduced communication). In 3D environments, each agent’s altitude/roll/pitch are locally measurable with respect to a common level-ground world frame (e.g., via altimeter or downward facing LiDAR and IMU). By constraining each agent to an altitude/roll/pitch envelope, this information only needs to be transmitted once a priori. These assumptions can then be locally monitored by each agent, with event-based communication only occurring when an agent changes or violates its constraint envelope (Figure 3(b)). This formulation allows us to simplify the 6-DoF optimization (i.e., 3D position and orientation) to a 3-DoF optimization (i.e., $x$, $y$, yaw with known $z$, roll, pitch) that, with fewer free variables, more reliably produces 3D localization solutions. Thus our approach provides a 3D relative pose using only locally collected UWB range measurements, a priori state constraints, and detections of constraint violations.
@@
This paper’s contributions are: (1) An in-depth analysis and modeling of the observed noise characteristics of 3D UWB ranging measurements (Section III). (2) A 3D instantaneous solution for UWB-based relative localization with minimal communication (i.e., only requires event-based communication when a locally monitored assumption is violated, which should occur infrequently or never). Our solution is formulated as a robust nonlinear least squares optimization with a learned measurement bias correction as a function of relative elevation, improving our mean absolute position error by 19% (Section IV). (3) Hardware experimental results that demonstrate the merits of our proposed solution, a mean APE and AHE of 0.24m and 9.5 respectively without continuously transmitting measurements (Section V). (4) Our datasets, over 200 hours of pairwise UWB range measurements with ground truth.444 https://github.com/mit-acl/murp-datasets
@@
Figure 2: Plots demonstrating the UWB noise and geometry characteristics outlined in Section III. (a) Histogram of our entire set of UWB measurements binned by range error. Demonstrates a non-zero mean and long tail (i.e., violates the zero mean Gaussian assumption that is typically used). (b) Same data plotted as error with respect to relative elevation showing that the measurement error’s mean and variance change significantly with relative elevation. The dotted blue line represents a learned 6-degree polynomial fit of measurement bias. (c) Simple example demonstrating dilution of precision (DOP) in a 2D ranging scenario. (d) Demonstrates how three (or more) ranging antennas within a single plane produce a pair of ambiguous solutions (i.e., if all antennas are in the base’s plane $z=0$, while the target’s $x$ and $y$ coordinates are fully observable, the target’s altitude has an ambiguity between ±$\pm z$).
@@ II Related Works
2D Guo et al.     (2020) [11] Single Bounded Error Assumption UWB + Velocity
Cao et al.     (2021) [7] Many (4) NLLS UWB + Odometry
Zheng et al.     (2022) [12] Many (2 or 4) ZM Gaussian + NLOS Rejection UWB + VIO
Zhang et al.     (2023) [13] Many (2 or 4) ZM Gaussian + Outlier Rejection UWB + Point Clouds + Odom + Keyframes
[Our Prev. Work]     (2022) [10] Many (4) NLLS + Data Informed Sensor Model UWB
3D Xu et al.     (2020) [8] Single ZM Gaussian UWB + VIO + Visual Tracks
Qi et al.       (2020) [14] Single ZM Gaussian UWB + IMU + GPS Heading
Xianjia et al.     (2021) [15] Many (2 or 4) NLLS UWB
Xu et al.     (2022) [9] Single ZM Gaussian + Occlusion Rejection UWB + VIO + Visual Tracks
Xun et al.     (2023) [16] Single ZM Gaussian UWB + IMU + IR Visual Tracks
[This Work]     (2024) Many (4 or 6) NLLS + Data Informed Sensor Model UWB + Event-Based Assumption Violations
@@
Zhang et al. [13] presents a centralized 2D range-aided cooperative localization and consistent reconstruction system, merging a tightly coupled visual odometry and multi-tag ranging front-end with a pose graph optimization (PGO) back-end. This work differs from ours by being 2D and relying on continuous transmission of point cloud, odometry, and keyframe information. Omni-swarm [9] is a decentralized 3D swarm estimation system. This work differs from ours by relying primarily on visual tracks of neighboring agents from each agent’s omnidirectional camera – omitting UWB measurements only degrades their estimation by approximately 0.01m. Additionally, each agent has only a single UWB tag and requires continuously sharing measurements between all pairs of agents, scaling poorly with larger swarms. CREPES [16] presents a custom compact hardware module, which tightly couples an IMU, UWB, IR LEDs, and IR fish-eye camera. The measurements are fused into 6-DoF relative pose via a centralized error-state Kalman filter (ESKF) and PGO, while also having the ability to provide an instantaneous estimate from a single frame of sensor measurements. This work differs from ours by relying heavily on visual tracks over UWB (similar to [9]). Additionally, the system relies on continuously transmitted measurements. Thus the state of the art approaches in [13, 9, 16] achieve APE and AHE on the order of 0.10m and $1^{\circ}$ respectively, but require additional sensors that must continuously transmit measurements, making them scale poorly to scenarios with larger swarm sizes or reduced communication throughput. Furthermore, these required additional sensors (e.g., LiDAR, specialized cameras, etc.) can cost thousands of dollars per agent, whereas UWB sensors (even multiple per robot) can cost $<$\$100 per agent.
@@
TABLE II: Theoretical standard deviation of absolute position error between two of our UAVs (illustrated in Figure 1 and Section V-A). Computation uses a UWB ranging uncertainty of $\sigma=0.24$m (see Figure 2(a)). Altitude assumptions are the standard deviation of the target’s assumed altitude when local only, and ±$\pm 0.04$m (the advertised precision of the downward LiDAR) when continuously sharing measurements. Since the system is nonlinear, table columns show evaluation at several relative altitudes. Table rows show how different relative altitude assumptions affect uncertainty. Results demonstrate the advantage of relative altitude assumptions and shows how said assumptions enable comparable uncertainty with only locally collected measurements vs systems that continuously share altitude measurements. See Section III-C for additional discussion.
[…]
StdDev of Abs Pos Error [m] where $\Delta z$ equals:
name local only alt. assumpt. 0.0m 1.0m 2.5m 5.0m 10.0m 25.0m
no_assumpt n/a $\infty$ 9.98 2.05 1.35 2.81 14.38
extr_conserv ±$\pm 2.00$m 4.02 2.89 1.43 1.21 2.66 13.73
very_conserv ±$\pm 1.00$m 1.08 1.00 0.84 1.00 2.38 12.48
conserv ±$\pm 0.20$m 0.39 0.41 0.49 0.79 1.96 10.21
proposed ±$\pm 0.10$m 0.39 0.41 0.49 0.78 1.96 10.17
cont_shared ±$\pm 0.04$m 0.39 0.41 0.49 0.78 1.96 10.16
@@ III Extensions to 3D Environments
The following subsections motivate our hardware and algorithmic choices for operating in 3D (Section III-C). We first characterize the UWB ranging error of the Nooploop LinkTrack P-B [2] from a set of initial experiments [^5] While these effects should generalize to other UWB ranging devices, the “black box” nature of commercial sensors means the exact manifestations could differ between products. Most works in Table I use a Nooploop LinkTrack P model [2], except for [11] (PulsON 440) and [15] (unspecified). Thus, we expect similar noise characteristics among those related works. (Section III-A), and then provide the observability and error properties of trilateration-based localization (Section III-B).
@@ III-A Characterizing UWB Noise
Long Tail Distributions: Contrary to common noise assumptions in Table I, individual UWB ranging errors appear neither zero mean nor Gaussian. Instead, error distributions have long positive tails (see Figure 2(a)). A trend towards positive bias can be attributed to the many ways positive ranging error can be introduced to a UWB measurement (e.g., multipath or a change of propagation medium). Furthermore, installing SMA cables between RF devices and antennas adds consistent positive bias to all measurements.
@@ III-B Geometry of Ranging-based Pose Estimation
Dilution of Precision: Most commonly associated with GPS, dilution of precision (DOP) is a trilateration sensitivity analysis for quantifying how 1D ranging errors propagate to 2D/3D point estimation error [23]. Specifically, given trilateration’s inherent nonlinearity, even identically noisy 1D range measurements can produce significantly different 2D/3D positional uncertainties – this is entirely determined by the given base station and target geometry (Figure 2(c)). In general we note that increasing the “baseline” distance between base stations (i.e., the antenna separation on a single vehicle), or having a target closer to the base stations (i.e., decreasing distance between two vehicles), reduces trilateration’s 2D/3D uncertainty. DOP also accounts for the improved accuracy seen by [15] for configurations with larger baselines. Thus, the accuracy of trilateration systems with vastly different baselines or operating ranges can be misleading. See [12, 13] for a detailed discussion of DOP as it pertains to 2D UWB localization.
@@ III-C System Design
Based on the noise characteristics (Section III-A) and underlying geometry (Section III-B), we propose the system design shown in Figure 1 for our multirotor drones. Specifically, each agent is equipped with 6 ranging antennas (attached 0.31m from the center at the end of each propeller arm) to maximize the baselines and improve DOP geometry. Since there are 6 coplanar antennas, this is an over-constrained problem that still has a ±$\pm z$ ambiguity across the $z=0$ plane. Although an additional nonplanar antenna could be added to each agent, it is challenging to get a large $z$-baseline without impacting the flight characteristics (i.e., heavy/awkward configuration). Instead, we note that altitude/roll/pitch can be measured within the flat-ground world frame locally via altimeter or downward facing LiDAR and onboard IMU. [^6] Altitude measurements are direct and should not drift. Substantial drift in roll/pitch w.r.t. gravity would impede basic flight operations before meaningfully impacting our estimation, making it negligible for our system. In many drone applications, it is common for multirotors to maintain roll/pitch near $0^{\circ}$ and a constant altitude. Thus, we can specify a minimalist communication protocol that shares these intended constraints a priori and then monitors them locally to ensure they are satisfied. Event-based communication would then only occur if a constraint is locally detected to have been violated/changed (Figure 3(b)).
Table II summarizes the theoretical uncertainty of the proposed system and demonstrates the value of adding relative altitude assumptions to a communication constrained system. [^7] See explanation and code in the GitHub repository for more information on how the theoretical values in Table II were computed. Although the assumed altitude bounds should vary with the platform, we see even overly conservative assumptions (e.g., ±$\pm 2.0$m) still greatly improves performance over no assumption (no_assumpt vs extr_conserv or very_conserv). This is especially true for level or near-level flight formations (i.e., unobservablility at $\Delta z=0$ and high sensitivity nearby), a common configuration for many UAV systems. By comparing the positional uncertainty of our proposed assumption (i.e., proposed, experimentally selected based on Figure 4(a)), a twice as lenient assumption (i.e., conserv), or continuously communicated altitude measurements (i.e., cont_shared, using the advertised downward LiDAR accuracy of ±$\pm 0.04$m) we see our minimal communication protocol does not substantially degrade uncertainty while using only locally collected measurements.
@@
Figure 3: (a) Annotated diagram of three UGV agents used in Section V-B. UGV agents are designed with the same baseline as the UAV in Figure 1, making them comparable surrogates to the UAV and its experiments in Section V-A. (b) Diagram demonstrating agent $B$ locally detecting violations in its a priori constraints, triggering an event-based communication. Other agents will exclude $B$ from their relative pose estimation until $B$ notifies the swarm of restored constraint status or provides a new constraint envelop.
IV Technical Approach